OpenStudy (abhisar):

Can anyone teach me how to write configuration for hetero-atomic molecules using MOT. Eg. \(\sf H_2O\)

OpenStudy (abhisar):

@Kainui

OpenStudy (abhisar):

@JFraser

OpenStudy (abhisar):

I am doing it like \(\sigma 1s^2 ~\sigma^*1s^2~\sigma 2s^2~\sigma ^*2s^2 (\pi 2p_x^1 \pi2p_y^1)\)

OpenStudy (abhisar):

But according to this there are 2 unpaired electrons and thus water should be para-magnetic but water is actually diamagnetic.

OpenStudy (abhisar):

@aaronq

OpenStudy (anonymous):

Your problem is here in the \(\sigma ^*2s\) orbital's placement. It's actually higher in energy than the \(\pi\) orbitals. \(\sigma 1s^2 ~\sigma^*1s^2~\sigma 2s^2~\sigma ^*2s^2 (\pi 2p_x^1 \pi2p_y^1)\) So the correct placement should be: \(\sigma 1s^2 ~\sigma^*1s^2~\sigma 2p_z^2(\pi 2p_x^2 \pi2p_y^2)~\sigma ^*2p_z^0 \) Which has only paired electrons, thus diamagnetic like you're looking for. Also note that I changed them from s to \(p_z\). Remember these \(\sigma\) molecular orbitals came from two p orbitals not s orbitals! Here's the difference in a picture!